OpenStudy (anonymous):

differentiate by rule: 3x-4/2x+5

OpenStudy (sidsiddhartha):

by quotient rule?

OpenStudy (sidsiddhartha):

\[\frac{ d }{ dx }[\frac{ f(x) }{ g(x) }]=\frac{ g(x).f'(x)-f(x).g'(x) }{ g(x)^2 }\]

OpenStudy (anonymous):

how did u do that

OpenStudy (sidsiddhartha):

its just the rule so here u have f(x)=3x-4 g(x)=2x+5 try with this